OpenStudy (anonymous):

You invest $300 in an account that has an interest rate of 1.3%, compounded monthly. How much money is in the account after 10 years? Round your answer to the nearest whole number. Answer $342 $345 $356 $360

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

i have forgotten the formula :(

OpenStudy (anonymous):

@ryan123345

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

A = P(1 + r/n)^(n*t) is the formula you will use

OpenStudy (anonymous):

@ghazi In case you forgot @jim_thompson5910 brought it up thanks I shall try it

OpenStudy (ghazi):

yes give a try

OpenStudy (anonymous):

I got 356? O.o

jimthompson5910 (jim_thompson5910):

that's incorrect

OpenStudy (anonymous):

either that or 345

jimthompson5910 (jim_thompson5910):

A = P(1 + r/n)^(n*t) A = 300(1 + 0.013/12)^(12*10) A = ???

OpenStudy (anonymous):

341.624

OpenStudy (anonymous):

so 342 is the answer.

jimthompson5910 (jim_thompson5910):

yep